OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet?

OpenStudy (anonymous):

x=width of box. length of box = 2 * x. Height of box = x - 3. Volume of box = length * width * height = 2x * x * (x-3) = 2x^2 * (x-3) = 2x^3 - 6x^2

OpenStudy (anonymous):

ok thanx alot

OpenStudy (anonymous):

welcome